OpenStudy (anonymous):

Ok if i'm 2.39 meters away from the base of a pole and to the top of it is a 55 degree angle from where i'm standing. How do I find the Height? I know i can put in (2)(tan(55) into my ti-83 but i was hoping that someone could break down the math for me so I can learn it.

OpenStudy (dumbcow):

 tan x = opp/adj tan(55) = h/2.39 h = 2.39*tan(55)